invariance under positive scaling W exp1 p X

  • Slides: 63
Download presentation

指數隨機變數的不變性 (invariance under positive scaling) 若W ∼ exp(1) p 且 X = θW, θ

指數隨機變數的不變性 (invariance under positive scaling) 若W ∼ exp(1) p 且 X = θW, θ > 0, p則 p X ∼ exp(θ) E(X)=E(θW)= θ E(1)= θ V(X)=V(θW)= θ 2 E(1)= θ 2

Exponential Distribution Example: Customers arrive at the service counter at the rate of 15

Exponential Distribution Example: Customers arrive at the service counter at the rate of 15 per hour. What is the probability that the arrival time between consecutive customers is less than three minutes? § § The mean number of arrivals per hour is 15, so § λ = 1/θ=60/4 =15 Three minutes is 0. 05 hours P(arrival time <. 05) = 1 – e-λX = 1 – e-(15)(0. 05) = 0. 5276 So there is a 52. 76% probability that the arrival time between successive customers is less than three minutes

Example p 1. 2. Suppose the earthquakes occur in the eastern part of Taiwan

Example p 1. 2. Suppose the earthquakes occur in the eastern part of Taiwan un accordance with the assumptions of the Poisson probability distribution at a rate of 2 per day. Find the probability that at least 3 earthquakes occur during the next 2 days. Find the probability distribution of the time, starting from now , until the next earthquake.

Johann Carl Friedrich Gauss (1777– 1855)

Johann Carl Friedrich Gauss (1777– 1855)

Normal Distribution Probability is area under curve! f(x) c d x

Normal Distribution Probability is area under curve! f(x) c d x

The Standard Normal Table: P(0 < z < 1. 96) Standardized Normal Probability Table

The Standard Normal Table: P(0 < z < 1. 96) Standardized Normal Probability Table (Portion) Z . 04 . 05 . 06 =1 1. 8. 4671. 4678. 4686 . 4750 1. 9. 4738. 4744. 4750 2. 0. 4793. 4798. 4803 = 0 1. 96 Z 2. 1. 4838. 4842. 4846 Probabilities Shaded area exaggerated

The Standard Normal Table: P(– 1. 26 z 1. 26) Standardized Normal Distribution =1.

The Standard Normal Table: P(– 1. 26 z 1. 26) Standardized Normal Distribution =1. 3962 P(– 1. 26 ≤ z ≤ 1. 26) =. 3962 +. 3962 – 1. 26 Z =0 Shaded area exaggerated =. 7924

The Standard Normal Table: P(z > 1. 26) Standardized Normal Distribution s=1 P(z >

The Standard Normal Table: P(z > 1. 26) Standardized Normal Distribution s=1 P(z > 1. 26) . 5000 =. 5000 –. 3962 1. 26 m=0 Z =. 1038

The Standard Normal Table: P(– 2. 78 z – 2. 00) Standardized Normal Distribution

The Standard Normal Table: P(– 2. 78 z – 2. 00) Standardized Normal Distribution =1 P(– 2. 78 ≤ z ≤ – 2. 00). 4973 =. 4973 –. 4772 =. 0201 . 4772 – 2. 78 – 2. 00 =0 Shaded area exaggerated Z

The Standard Normal Table: P(z > – 2. 13) Standardized Normal Distribution =1. 4834

The Standard Normal Table: P(z > – 2. 13) Standardized Normal Distribution =1. 4834 P(z > – 2. 13) . 5000 =. 4834 +. 5000 – 2. 13 =0 Shaded area exaggerated Z =. 9834

Standardize the Normal Distribution Standardized Normal Distribution = 1 X = 0 One table!

Standardize the Normal Distribution Standardized Normal Distribution = 1 X = 0 One table! Z

Non-standard Normal μ = 5, σ = 10: P(5 < X< 6. 2) Normal

Non-standard Normal μ = 5, σ = 10: P(5 < X< 6. 2) Normal Distribution Standardized Normal Distribution = 10 =1. 0478 = 5 6. 2 X Shaded area exaggerated = 0. 12 Z

Non-standard Normal μ = 5, σ = 10: P(3. 8 X 5) Normal Distribution

Non-standard Normal μ = 5, σ = 10: P(3. 8 X 5) Normal Distribution Standardized Normal Distribution =1 = 10 . 0478 3. 8 = 5 X Shaded area exaggerated -. 12 = 0 Z

Non-standard Normal μ = 5, σ = 10: P(2. 9 X 7. 1) Normal

Non-standard Normal μ = 5, σ = 10: P(2. 9 X 7. 1) Normal Distribution Standardized Normal Distribution = 10 =1. 1664. 0832 2. 9 5 7. 1 X Shaded area exaggerated -. 21 0. 21 Z

Non-standard Normal μ = 5, σ = 10: P(X 8) Normal Distribution Standardized Normal

Non-standard Normal μ = 5, σ = 10: P(X 8) Normal Distribution Standardized Normal Distribution = 10 =1. 5000. 1179 =5 8 X Shaded area exaggerated =0 . 3821 . 30 Z

Non-standard Normal μ = 5, σ = 10: P(7. 1 X 8) Normal Distribution

Non-standard Normal μ = 5, σ = 10: P(7. 1 X 8) Normal Distribution Standardized Normal Distribution = 10 =1. 1179. 0832 =5 7. 1 8 X =0 . 21. 30 . 0347 Z

Normal Distribution Thinking Challenge You work in Quality Control for GE. Light bulb life

Normal Distribution Thinking Challenge You work in Quality Control for GE. Light bulb life has a normal distribution with = 2000 hours and = 200 hours. What’s the probability that a bulb will last A. between 2000 and 2400 hours? B. less than 1470 hours?

Solution* P(2000 X 2400) Normal Distribution Standardized Normal Distribution = 200 =1 . 4772

Solution* P(2000 X 2400) Normal Distribution Standardized Normal Distribution = 200 =1 . 4772 = 2000 2400 X =0 2. 0 Z

Solution* P(X 1470) Normal Distribution Standardized Normal Distribution = 200 =1. 5000 . 0040

Solution* P(X 1470) Normal Distribution Standardized Normal Distribution = 200 =1. 5000 . 0040 1470 = 2000 X . 4960 -2. 65 = 0 Z

Finding Z Values for Known Probabilities Standardized Normal Probability Table (Portion) What is Z,

Finding Z Values for Known Probabilities Standardized Normal Probability Table (Portion) What is Z, given P(Z) =. 1217? . 1217 =1 Z . 00 . 01 0. 2 0. 0. 0000. 0040. 0080 0. 1. 0398. 0438. 0478 =0 Shaded area exaggerated . 31 ? Z 0. 2. 0793. 0832. 0871 0. 3. 1179. 1217. 1255

Empirical Rules What can we say about the distribution of values around the mean?

Empirical Rules What can we say about the distribution of values around the mean? For any normal distribution: f(X) σ μ-1σ μ ± 1σ encloses about 68. 26% of X’s σ μ μ+1σ 68. 26% X

The Empirical Rule p μ ± 2σ covers about 95% of X’s p μ

The Empirical Rule p μ ± 2σ covers about 95% of X’s p μ ± 3σ covers about 99. 7% of X’s 2σ 3σ 2σ μ 95. 44% x 3σ μ 99. 73% x